### v3

# On Rate-Optimal Partitioning Classification from Observable and from Privatised Data

Abstract
In this paper we revisit the classical method of partitioning classification and study its convergence rate under relaxed conditions, both for observable (non-privatised) and for privatised data. We consider the problem of classification in a $d$ dimensional Euclidean space. Previous results on the partitioning classifier worked with the strong density assumption, which is restrictive, as we demonstrate through simple examples. Here, we study the problem under much milder assumptions. We presuppose that the distribution of the inputs is a mixture of an absolutely continuous and a discrete distribution, such that the absolutely continuous component is concentrated to a $d_{a}$ dimensional subspace. In addition to the standard Lipschitz and margin conditions, a novel characteristic of the absolutely continuous component is introduced, by which the exact convergence rate of the classification error probability is computed, both for the binary and for the multi-label cases. Interestingly, this rate of convergence depends only on the intrinsic dimension of the inputs, $d_{a}$. The privacy constraints mean that the independent identically distributed data cannot be directly observed, and the classifiers are functions of the randomised outcome of a suitable local differential privacy mechanism. In this paper we add Laplace distributed noises to the discontinuations of all possible locations of the feature vector and to its label. Again, tight upper bounds on the rate of convergence of the classification error probability are derived, without the strong density assumption, such that this rate depends on $2d_{a}$.
62G05,
62G20,
classification,
local differential privacy,
partitioning rule,
rate of convergence,
keywords:
[class=AMS]
keywords:
[…]
, , and
1 Introduction
Classification is one of the fundamental problems of statistics and machine learning, see Devroye, Györfi and Lugosi [6] and Vapnik [20]. It has countless applications from health care and industry to commerce and finance. A significant portion of these applications include sensitive data, for example, about the health or financial situations of the clients involved, which should be anonymised before it can be processed. Nevertheless, anonymisation is more involved than just removing the names and the addresses of the clients, as the data can contain several other types of sensitive information. The concept of differential privacy in Dwork et al. [8] provides a rigorous framework to measure the amount of information privacy. Localised privacy dates back to Warner [21], and local differential privacy (LDP) was formally defined by Duchi, Jordan, and Wainwright [7]. On the one hand, LDP helps getting rid of a trusted third party; on the other hand, it generates a framework for managing nonparamteric regression and classification, cf. Berrett and Butucea [3] and Berrett, Györfi and Walk [4]. The LDP mechanism allows processing data even in cases, when the original, raw data should only be seen by its legitimate data holder.
In this paper, we focus on partitioning rules, see Kohler and Krzyżak [14], which are archetypical classification methods. They are key componenents of several machine learning methods, such as decision trees, random forrests, piecewise estimators and hierarchical models. They provide viable solutions to challenges related to scalability, interpretability and explainability, henceforth, they are well-suited for integration into, e.g., federated learning [1] and ensemble-based methods [5]. They are especially useful in low-dimensional settings, provide benchmark models for complex data structures and can be applied for hyperparameter tuning and data cleaning. In general, local averaging estimators are easier to understand compared to complex models and can be used to prove rate of convergence and minimax results in nonparametric settings. The main advantages of the partitioning classifier and its generalizations, like prototype classifiers are their small computational complexity and the sample compression property. Our aim is to analyse the convergence rate of its classification error probability under mild statistical assumptions. We investigate both binary and multi-label classification, covering observable and privatised (anonymised) data. In the latter case, we apply Laplace type randomisation which ensures LDP constraints, cf. [3].
Partitioning classification is often studied under Lipschitz and margin types conditions, but the existing optimal error bounds (for the non-privatised case) suppose a further condition, called the strong density assumption (SDA). This assumption, stated in (4), ensures that the probability measure of each cell of the partition is bounded away from zero. For observable (non-privatised) data Kohler and Krzyżak [14] studied the convergence rate of plug-in partitioning classification with and without SDA. Their rate with SDA was proven to be miminax optimal by Audibert and Tsybakov [2]. For the privatised case Berrett, Györfi and Walk [4] conjectured that without such an assumption, the convergence rate could be arbitrarily slow.
Nevertheless, SDA is fairly restrictive as we demonstrate through Examples 1, 2 and 3, for which SDA is not satisfied. In general, the Lipschitz and the margin parameters do not determine the convergence rate of the error probability. While the approximation error (bias) depends only on the Lipschitz and margin parameters, the estimation error is very sensitive to the behaviour of the distribution of the feature vector around the decision boundary. These motivate the introduction of an additional parameter characterizing the relation between the margin and the low density areas. Specifically, we assume that the distribution of input $X$ is a mixture of an absolutely continuous and a discrete distribution, such that the absolutely continuous component is concentrated on a $d_{a}$-dimensional subspace. For the absolutely continuous component, we introduce an additional mild condition, called the combined margin and density assumption, for which the true convergence rate of the estimation error can be derived. The combined margin and density parameter is an appropriate characterisation of the intersection of the low density region and the decision boundary.
Our convergence rate for the binary case, proved without SDA, greatly improves the SDA independent rate of Kohler and Krzyżak [14], see (5), as it is demonstrated on our specific examples with appropriately chosen parameters. Our result ensures the same optimal rate for Example 1 as the SDA dependent rate of Kohler and Krzyżak, that is (6). Hence, the optimal rate can be achieved even without SDA. Moreover, for Example 2 our bound matches the convergence rate one gets by direct calculation of the rate, showing that the bound is tight. Finally, for Example 3, though the deduced SDA independent convergence rate is worse than the SDA dependent rate of (6), but it still improves the previously known SDA independent rate, that is (5). Furthermore, interestingly, this convergence rate only depends on the intrinsic dimension of the absolutely continuous part, $d_{a}$. We also derive upper bounds for the convergence rate for the case of privatised data with LDP guarantees, which rate depends on $2d_{a}$, instead of $d_{a}$ that was the case for the nonprivate bound. Both of these bounds are extended to multi-label classification, as well. Our bounds for privatised partitioning classification refute the conjecture that without SDA the convergence rate could be arbitrarily slow.
The structure of the paper is as follows. First, in Section 2, we revisit and improve the error probability bounds of partitioning classifiers for observable data, both for binary and multi-labeled setups. Then, in Section 3, we follow an analogous program for privatised partitioning classifiers. The results are summarized and discussed in Section 4 and, finally, the detailed proofs are presented in Section 5.
2 Partitioning classification from observable data
In this section, we study the problem of classification from observable (non-privatised) data. First, we give an overview about the core problem, recall the partitioning classification rule, and state our main assumptions under which we quantify the convergence rate both for binary and for multi-label classification.
The standard setup of binary classification is as follows: let the random feature vector $X$ take values in ${\mathbb{R}}^{d}$, and let its label $Y$ be $\pm 1$ valued. We denote by $\mu$ the distribution of $X$, that is, $\mu(A)={\mathbb{P}}(X\in A)$ for all measurable sets $A\subseteq{\mathbb{R}}^{d}$. The task of classification is to decide on $Y$ given $X$, i.e., one aims to find a decision function $D$ defined on the range of $X$ such that $D(X)=Y$ with large probability. If $D$ is an arbitrary (measurable) decision function, then its error probability is denoted by
$L(D)={\mathbb{P}}\{D(X)\neq Y\}.$
It is well-known that the Bayes decision $D^{*}$ minimizes the error probability:
$D^{*}(x)=\mbox{sign}(m(x)),$
where $\mbox{sign}(x)={\mathbb{I}}_{\{x\geq 0\}}-{\mathbb{I}}_{\{x<0\}}$, with the regression function
$m(x)={\mathbb{E}}[Y|X=x],$
which is well-defined for $\mu$-almost all $x\in{\mathbb{R}}^{d}$. Then,
$L^{*}={\mathbb{P}}\{D^{*}(X)\neq Y\}=\min_{D}L(D)$
denotes its error probability (i.e., the probability of misclassification).
Let ${\cal P}_{h}=\{A_{h,1},A_{h,2},\ldots\}$ be a cubic partition of ${\mathbb{R}}^{d}$ with cubic cells $A_{h,j}$ of volume $h^{d}$ such that $(0,h]^{d}\in{\cal P}_{h}$. The raw data $\mathcal{D}_{n}$ are assumed to be independent identically distributed (i.i.d.) copies of the random vector $(X,Y)$,
$\mathcal{D}_{n}=\{(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})\}.$ (1)
Put
$\nu_{n}(A_{h,j})=\frac{1}{n}\sum_{i=1}^{n}Y_{i}\,{\mathbb{I}}_{\{X_{i}\in A_{h,j}\}},$
where $\mathbb{I}$} stands for the indicator function. The well-known partitioning classification rule is
$D_{n}(x)=\mbox{sign}(\nu_{n}(A_{h,j})),\quad\mbox{if}\;\;x\in A_{h,j}.$
A nontrivial rate of convergence of any classification rule can be derived under some smoothness condition. The Lipschitz condition on $m$ means that there is a constant $C$ such that for all $x,z\in{\mathbb{R}}^{d}$, we have
$|m(x)-m(z)|\leq C\,\|x-z\|.$ (2)
Concerning the rate of convergence of any classification rule, Mammen and Tsybakov [15] and Tsybakov [17] discovered and investigated the phenomenon that there is a dependence on the behaviour of $m$ in the neighborhood of the decision boundary
$B^{*}=\{x:m(x)=0\}.$
The margin condition means that for all $0<t\leq 1$, we have
$G^{*}(t)$ $:=\int{\mathbb{I}}_{\{0<|m(x)|\leq t\}}\mu(dx)\leq c^{*}\cdot t^{\gamma},$ (3)
for some constants $c^{*}\geq 0$ and $\gamma\geq 0$.
We say that the strong density assumption (SDA) holds, when for all $\mu(A_{h,j})>0$, we have
$\mu(A_{h,j})\geq ch^{d},\qquad j=1,\ldots,$ (4)
for some constant $c>0$. SDA is an extremely restrictive condition, because, for example, it excludes the continuous densities supported on ${\mathbb{R}}^{d}$, including the case of the classical discriminant analysis, when the conditional distributions of $X$ given $Y$ are Gaussian.
If $X$ is bounded, and the margin and the Lipschitz conditions on $m$ are satisfied, then Kohler and Krzyżak [14] showed for partitioning classification with suitably chosen $(h_{n})$ that
${\mathbb{E}}\{L(D_{n})\}-L^{*}$ $=O\left(n^{-\frac{1+\gamma}{3+\gamma+d}}\right).$ (5)
If, in addition, the SDA is met, then the order of the rate of convergence is
$n^{-\frac{1+\gamma}{2+d}}.$ (6)
Under SDA, Audibert and Tsybakov [2] proved the minimax optimality of this rate. In general, the Lipschitz and the margin conditions do not determine the true rate of convergence of the error probability.
Now, we consider three examples, when the SDA is not satisfied.
Example 1.
Let the range of $X$ be the interval [1,1][-1,1] and
$m(x)=x,\quad|x|\leq 1.$
Furthermore, assume that $\mu$ has the density
$f(x)=c_{\delta}(1-|x|^{\delta}),\quad|x|<1,$
with $\delta>0$ and $c_{\delta}$ being a normalizing constant.
Example 2.
Let $m$ be as in Example 1. Assume that $\mu$ has the density
$f(x)=c_{\delta}|x|^{\delta},\quad 0<|x|\leq 1,$
with $\delta>-1$ and normalizing constant $c_{\delta}$.
Example 3.
Let the range of $X$ be the interval [1,1][-1,1] and
$m(x)=\mbox{sign }(x)\cdot x^{2},\quad|x|\leq 1.$
Assume that $\mu$ has the density
$f(x)=|x|,\quad|x|\leq 1.$
In Example 1 the margin condition holds with $\gamma=1$. Thus, in the literature only the poor rate (5) has been proved, which is $n^{-2/5}$. By an easy calculation, for the choice $h_{n}=n^{-1/3}$, we obtain that the true rate is $n^{-2/3}$, which corresponds to the optimal rate (6). This means that there is space for improvement, i.e., we can improve the rate if the SDA is weakened. The interesting feature of Example 2 is that the density tends to infinity at the decision boundary for $1<\delta<0$, and the rate of convergence is worse than $n^{-2/3}$, see the details after Theorem 2.1. In Example 3 the density tends to zero at the decision boundary for $m$.
Motivated by these examples, we introduce an additional parameter characterizing the relation between the margin and the low density areas. We assume a decomposition
$\mu=\mu_{a}+\mu_{s},$ (7)
where $\mu_{a}$ is an arbitrary absolutely continuous distribution on its support $S_{a}$, which is contained in an unknown $d_{a}$ dimensional Euclidean space, and let $f$ be its density with respect to the $d_{a}$ dimensional Lebesgue measure $\lambda_{d_{a}}$. In this paper the density $f$ is arbitrary, i.e., we don’t assume any regularity property on $f$. Furthermore, assume that $\mu_{s}$ is a discrete distribution with support $S_{s}$ of finite size. For this setup, our motivation was the example when $X$ has $d_{a}$ continuous features and it has $d_{s}$ discrete features.
Under the Lipschitz and margin condition, one can derive tight bounds on the approximation error component of the classifier’s error probability. The novel definition below is on the relation between the margin area and the low density region. This concept will be a useful tool in the refined analysis of the rate of convergence of the estimation error.
Put
$f_{h}(x)=\frac{\mu_{a}(A_{h,j})}{\lambda_{d_{a}}(A_{h,j})}=\frac{\int_{A_{h,j}}f(z)\lambda_{d_{a}}(dz)}{\lambda_{d_{a}}(A_{h,j})},\quad\mbox{ if }x\in A_{h,j}.\vskip 2.84526pt$ (8)
One can observe that function $f_{h}$ is the expectation of the histogram density estimate.
The combined margin and density condition intuitively controls the measure of a region which is either close to the decision boundary or is such that the density of $X$ is small. Assume that there exists $h_{1}^{*}>0$ such that for any $h\in(0,h_{1}^{*})$ and for all $t>0,$ we have
$G_{h}(t):=\int_{S_{a}}{\mathbb{I}}_{\{0<\sqrt{f_{h}(x)}|m(x)|\leq t\}}\frac{f(x)}{\sqrt{f_{h}(x)}}\,\lambda_{d_{a}}\!(dx)$ $\leq c_{1}\cdot t^{\gamma_{1}}$ (9)
with constants $c_{1}>0$ and $0\leq\gamma_{1}=\gamma_{1}(\sqrt{f})$.
The generalized Lebesgue density theorem yields that for $\lambda_{d_{a}}$-almost all $x$,
$\lim_{h\downarrow 0}f_{h}(x)$ $=f(x),$ (10)
cf. Theorem 7.2 in Wheeden and Zygmund [22]. Because of (10), we conjecture that if
$\int_{S_{a}}{\mathbb{I}}_{\{0<\sqrt{f(x)}|m(x)|\leq t\}}\frac{1}{\sqrt{f(x)}}\mu_{a}(dx)$ $\leq\tilde{c}\cdot t^{\gamma_{1}}$
with a $\tilde{c}>0$ and $\gamma_{1}\geq 0$, then the combined margin and density condition holds.
Under mild regularity conditions, the following theorem contains tight upper bounds on the rate of convergence of the error probability. Interestingly, only the dimension $d_{a}$ of the absolutely continuous component matters. The proof is presented in Section 5.
Theorem 2.1.
Assume that $X$ is bounded, $m$ satisfies the Lipschitz condition (2) and the margin condition (3) with $0\leq\gamma$. In addition, the combined margin and density condition (9) holds with $\gamma_{1}\geq 0$. Let $(h_{n})$ be a monotonic decreasing null sequence. Then, we have
$L(D_{n})-L^{*}$ $=O\left(h_{n}^{1+\gamma}\right)+O\left((nh_{n}^{d_{a}})^{-(1+\min\{1,\gamma,\gamma_{1}\})/2}\right).$ (11)
For known $d_{a}$ and for the choice
$h_{n}$ $=n^{-\frac{1}{2+d_{a}}},$ (12)
one has that
${\mathbb{E}}\{L(D_{n})\}-L^{*}$ $=O\Big{(}n^{-\frac{1+\min\{1,\gamma,\gamma_{1}\}}{2+d_{a}}}\Big{)}.$ (13)
If $d_{a}$ is not known, then choosing $h_{n}=n^{-\frac{1}{2+d}}$ yields a slightly worse bound, where the first term is $O\big{(}n^{-\frac{1+\gamma}{2+d}}\big{)}$.
In Example 1 for all $\delta>0$, the margin condition and the combined margin and density condition hold such that $\gamma=1$ and $\gamma\leq\gamma_{1}$ because of Lemma 2.2 below. Thus, (13) results in the rate $n^{-2/3}$, which is the same as the optimal rate in (6). In Example 2, one has $\gamma=\delta+1$ and $\gamma_{1}\geq 1$, therefore the bound on the rate in (13) is equal to $n^{-(2+\delta)/3}$ for $1<\delta\leq 0$ and $n^{-2/3}$ for $\delta>0$. For the choice $h_{n}=n^{-1/3}$, by direct calculation of the error probability, one gets the same convergence rate, hence the bound (13) is tight. In Example 3, one has $\gamma=1$ and $\gamma_{1}=3/5$ and so $\gamma>\gamma_{1}$. Thus, the rate in Theorem 2.1 is $n^{-8/15}$, which is faster than the poor rate in (5), but it is worse than the rate in (6) with $\gamma=1$.
The combined margin and density condition always holds with $\gamma_{1}=0$ in which case for the choice of (12), (13) says that
${\mathbb{E}}\{L(D_{n})\}-L^{*}$ $=O\Big{(}\,n^{-\frac{1}{2+d_{a}}}\,\Big{)}.$
In the bound (11) the first term, called approximation error bound, follows from the Lipschitz condition and from the margin condition, while the second term, called estimation error bound, has been derived from the margin condition and from the combined margin and density condition. Note that the upper bound on the estimation error is managed by CLT approximation with an error term of order $O(1/(nh_{n}^{d_{a}}))$. Therefore, the estimation error is upper bounded by $c_{1}\cdot(nh_{n}^{d_{a}})^{-(1+\min\{1,\gamma,\gamma_{1}\})/2}+c_{0}/(nh_{n}^{d_{a}})$, and it is lower bounded by $c^{\prime}_{1}\cdot(nh_{n}^{d_{a}})^{-(1+\min\{1,\gamma,\gamma_{1}\})/2}-c_{0}/(nh_{n}^{d_{a}})$.
The following lemma is on the particular setups, when SDA holds in the neighborhood of the decision boundary $B^{*}$.
Lemma 2.2.
For  $0<\epsilon<1$  and  $0<h_{0}$, set
$B^{*}_{\epsilon}=\{x:|m(x)|\leq\epsilon\},$
and $f_{\epsilon,min}=\inf_{x\in B^{*}_{\epsilon},0<h<h_{0}}f_{h}(x)$. If there is an $\epsilon\in(0,1)$ such that $f_{\epsilon,min}>0$, then $\gamma_{1}\geq\gamma$, therefore $\gamma=\min\{\gamma,\gamma_{1}\}$.
Next, we consider the classification problem for multiple classes. It means that $Y$ takes values in $\{1,\dots,M\}$. Let
$P_{k}(x)={\mathbb{P}}\{Y=k\mid X=x\}$
denote the a posteriori probabilities for $k=1,\dots,M$. Then, the Bayes decision has the form
$D^{*}(x)=\mathop{\rm arg\,max}_{k}P_{k}(x).$
Let $P_{n,k}$ be an estimate of $P_{k}$ based on ${\cal D}_{n}$, then the plug-in classification rule $g_{n}$ derived from $P_{n,k}$ is defined by
$D_{n}(x)=\mathop{\rm arg\,max}_{k}P_{n,k}(x).$
Recently, [19] and [16] generalized the margin condition to the multi-class setting: let $P_{(1)}(x)\geq\dots\geq P_{(M)}(x)$ be the ordered values of $P_{1}(x),\dots,P_{M}(x)$. For multiple classes, the margin condition means that there are some $\gamma\geq 0$ and $c^{*}\geq 0$ such that
$G^{*}(t):=\int{\mathbb{I}}_{\{0<P_{(1)}(x)-P_{(2)}(x)\leq t\}}\mu(dx)\leq c^{*}t^{\gamma},\qquad 0<t\leq 1.$ (14)
Using this concept of margin condition, Györfi and Weiss [13] calculated the rate of convergence of a nearest neighbor based prototype classifier, when the feature space is a separable metric space.
For multiple classes, the combined margin and density condition means that there is a $h_{1}^{*}>0$ such that for any $h\in(0,h_{1}^{*})$ and for all $0<t,$ we have
$G_{h}(t):=\int_{S_{a}}{\mathbb{I}}_{\{0<\sqrt{f_{h}(x)}(P_{(1)}(x)-P_{(2)}(x))\leq t\}}\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx)$ $\leq c_{1}\cdot t^{\gamma_{1}}$ (15)
with constants $0<c_{1}$ and $0\leq\gamma_{1}$.
For multiple classes, the partitioning rule is defined by
$\nu_{n,k}(A_{h,j})=\frac{1}{n}\sum_{i=1}^{n}{\mathbb{I}}_{\{Y_{i}=k,X_{i}\in A_{h,j}\}},$
and the corresponding plug-in rule as
$D_{n}(x)=\mathop{\rm arg\,max}_{k}\nu_{n,k}(A_{h,j}),\quad\mbox{ for }x\in A_{h,j}.$
Our main result for the multiclass problem is as follows. Its proof is presented in Section 5.
Theorem 2.3.
Assume that $X$ is bounded. Additionally, assume that $P_{1},\dots,P_{M}$ satisfy the Lipschitz condition (2), the margin condition (14) with $0\leq\gamma$ and the combined margin and density condition (15) with $\gamma_{1}\geq 0$. Let $(h_{n})$ be a monotonically decreasing sequence with zero limit. Then, we have
${\mathbb{E}}\{L(\widetilde{D}_{n})\}-L^{*}$ $=O\left(h_{n}^{1+\gamma}\right)+O\left((nh_{n}^{d_{a}})^{-(1+\min\{1,\gamma,\gamma_{1}\})/2}\right).$
3 Partitioning classification under local differential privacy
One of the main purposes of this paper is to bound the error probability of partitioning classifiers in the case, when the raw data $\mathcal{D}_{n}$ is not directly accessible, but only a suitably anonymised surrogate. More precisely, the anonymised data must satisfy a local differential privacy (LDP) condition [3, 7]. Our work is motivated by Berrett and Butucea’s recent paper [3], where the first step in this direction was done. Notice that the same privatization mechanism was studied for regression and density estimation in [12] and [11], respectively.
Let us now state the privacy mechanism that we consider in this work for the anonymisation of the raw data $\mathcal{D}_{n}$ in (1). Our approach follows the technique of Laplace perturbation already considered in Berrett and Butucea [3]. In our privacy setup, the dataholder of the $i$-th datum $X_{i}$ generates and transmits the data
$Z_{i,j}=Y_{i}{\mathbb{I}}_{\{X_{i}\in A_{h,j}\}}+\sigma_{Z}\epsilon_{i,j},\quad j=1,\ldots\vskip 1.42262pt$
to the statistician, where the noise level is $\sigma_{Z}>0$, and $\epsilon_{i,j}$ ($i=1,\ldots,n$, $j=1,\ldots$) being independent centered Laplace random variables with unit variance. This means that individual $i$ generates noisy data for any cell $A_{h,j}$.
Now, we briefly recall the definition of LDP. Non-interactive privacy mechanisms can be described by the conditional distributions $Q_{i}$ of the privatized data $Z_{i}$, for $i=1,\dots,n$, where each $Z_{i}$ takes its values from a measurable space $(\mathcal{Z},\mathscr{Z})$. Specifically, given a realization of the raw data $(X_{i},Y_{i})=(x_{i},y_{i})$, one generates $Z_{i}$ according to the probability measure defined by $Q_{i}(A\mid(X_{i},Y_{i})=(x_{i},y_{i}))$, for any $A\in\mathscr{Z}$. Such a non-interactive mechanism is local since any data holder can independently generate privatized data (e.g., without a trusted third party). For a privacy parameter $\alpha\in[0,\infty]$, a non-interactive privacy mechanism is said to be an $\alpha$-locally differentially private mechanism if the condition
$\frac{Q_{i}(A\mid(X_{i},Y_{i})=(x,y))}{Q_{i}(A\mid(X_{i},Y_{i})=(x^{\prime},y^{\prime}))}\leq\exp(\alpha)\vskip 2.84526pt$
holds for all $A\in\mathscr{Z}$ and all realizations $(x,y)$, $(x^{\prime},y^{\prime})$ of the raw data. The noise level $\sigma_{Z}$ has to be chosen of the form $2\sqrt{2}/\alpha$, to make the overall mechanism satisfy $\alpha$-LDP [3].
For privatised data, Berrett and Butucea [3] introduced the partitioning classification rule:
${\widetilde{\nu}}_{n}(A_{h,j})=\frac{1}{n}\sum_{i=1}^{n}Z_{i,j},$
and
$\widetilde{D}_{n}(x)=\mbox{sign }{\widetilde{\nu}}_{n}(A_{h,j}),\quad\mbox{ if }x\in A_{h,j}.$
If $h=h_{n}\to 0$ and $\alpha=\alpha_{n}\to 0$ such that $n\alpha_{n}^{2}h_{n}^{2d}\to\infty$, then Berrett and Butucea [3] proved the universal consistency of the partitioning classifier and calculated the minimax rate
$(n\alpha_{n}^{2})^{-\frac{1+\gamma}{2(1+d)}}$
in the class, when the margin condition and the Lipschitz condition together with SDA hold.
Let us introduce the modified combined margin and density condition for privatization. We say that $m$ satisfies the modified combined margin and density condition if there exists $h_{2}^{*}$ such that for all $h\in(0,h_{2}^{*})$ we have for all $t>0$:
$\widetilde{G}_{h}(t)\doteq\int{\mathbb{I}}_{\{0<f_{h}(x)(P_{(1)}(x)-P_{(2)}(x))\leq t\}}\frac{1}{f_{h}(x)}\mu_{a}(dx)\leq c_{2}\cdot t^{\gamma_{2}}$ (16)
with $c_{2}>0$ and $\gamma_{2}\geq 0$.
The next theorem is the extension of Theorem 2.1 to the setup of privatised data. The rate of convergence differs in two ways w.r.t. the previous result. The estimation error grows with $1/h^{-2d_{a}}$, instead of $1/h^{-d_{a}}$ which was the case for (11). Besides, the rate of convergence depends on $\gamma_{2}$, which is the modified combined margin and density condition parameter. The proof of Theorem 3.1 is included in Section 5.
Theorem 3.1.
If $X$ is bounded, $m$ satisfies the Lipschitz condition (2), the margin condition (3) with $\gamma\geq 0$ and the combined margin and density condition (9) with $\gamma_{1}$, the modified combined margin condition with $\gamma_{2}$. Let $(h_{n})$ be a monotonically decreasing sequence with zero limit, then we have
${\mathbb{E}}(L(\widetilde{D}_{n}))-L^{*}$ $=O\left(h_{n}^{1+\gamma}\right)+O\left(\bigg{(}\frac{1}{nh_{n}^{d_{a}}}\bigg{)}^{(\min\{1,\gamma,\gamma_{1})\}+1)/2}\right)$
$\quad+O\left(\left(\frac{\sigma_{Z}^{2}}{nh_{n}^{2d_{a}}}\right)^{(1+\min\{\gamma,\gamma_{2}\})/2}\right)+O\left(\frac{\sigma_{Z}}{nh_{n}^{d}}\right).$
If for all $n\in\mathbb{N}^{+}$:
$h_{n}\leq n^{\frac{\min(1,\gamma,\gamma_{1})-\min(\gamma,\gamma_{2})}{d_{a}(2\min(\gamma,\gamma_{2})+1-\min(1,\gamma,\gamma_{1}))}}$
and $d\leq 2d_{a}$, then the second and fourth terms are dominated by the third term, hence for
$h_{n}=(n/\sigma_{Z}^{2})^{-\frac{1}{2+2d_{a}}},$
one has
${\mathbb{E}}\{L(\widetilde{D}_{n})\}-L^{*}$ $=O\left((n/\sigma_{Z}^{2})^{-\frac{1+\min\{1,\gamma,\gamma_{2}\}}{2+2d_{a}}}\right).$
For non-binary classification from privatised data, we assume the data
$Z_{i,j,k}={\mathbb{I}}_{\{Y_{i}=k\}}{\mathbb{I}}_{\{X_{i}\in A_{h,j}\}}+\sigma_{Z}\epsilon_{i,j,k},\quad j=1,\ldots$
Set
${\widetilde{\nu}}_{n,k}(A_{h,j})=\frac{1}{n}\sum_{i=1}^{n}Z_{i,j,k},$
and
$\widetilde{D}_{n}(x)=\mathop{\rm arg\,max}_{k}{\widetilde{\nu}}_{n,k}(A_{h,j}),\quad\mbox{ if }x\in A_{h,j}.$
Our final theorem, which is proved in Section 5, is the multi-label version of Theorem 3.1.
Theorem 3.2.
If $X$ is bounded, $P_{1},\dots,P_{M}$ satisfy the Lipschitz condition (2), the margin condition (3) with $\gamma\geq 0$, the combined margin and density condition (9) with $\gamma_{1}$ and the modified combined margin condition with $\gamma_{2}$. Let $(h_{n})$ be a monotonically decreasing null sequence, then we have
${\mathbb{E}}\{L(\widetilde{D}_{n})\}-L^{*}$ $=O\left(h_{n}^{1+\gamma}\right)+O\left(\bigg{(}\frac{1}{nh_{n}^{d_{a}}}\bigg{)}^{(\min\{1,\gamma,\gamma_{1})\}+1)/2}\right)$
$\quad+O\left(\bigg{(}\frac{\sigma_{Z}^{2}}{nh_{n}^{2d_{a}}}\bigg{)}^{(1+\min\{\gamma,\gamma_{2}\})/2}\right)+O\left(\frac{\sigma_{Z}}{nh_{n}^{d}}\right).$
4 Discussion
In this paper we investigated both the binary and the multiclass versions of partitioning classification. Studying these methods also helps better understanding a wide range of local averaging estimators, and it is directly relevant for various statistical and machine learning methods, including federated learning and ensemble-based approaches. One of our main contributions was that we weakened the strong density assumption (SDA) and proved fast convergence rates under the margin condition and a newly introduced combined margin and density condition. It was shown that these convergence rates can be achieved even without using SDA, refuting the conjecture that without SDA the convergence rate can be arbitrarily slow.
We extended our results for the (locally differentially) private partitioning algorithm, which has no direct access to the data, and we proved convergence rates under the margin condition, combined margin and density condition and the modified combined margin and density condition. As expected, the convergence rates of the nonprivate algorithms are faster than the corresponding privatized ones. Our theorems quantify the effect of privatization by introducing an extra rate of $2$ for the bandwidth of the partitions and adjust the dependence of the rate to the density of the inputs by using a modified combined margin and density parameter.
Acknowledgements
The research of László Györfi was supported by the National Research, Development and Innovation Office (NKFIH) of Hungary under the 2023-1.1.1-PIACI-FÓKUSZ-2024-00051 funding scheme. The work of Balázs Csanád Csáji and Ambrus Tamás was supported, in part, by the European Union project RRF-2.3.1-21-2022-00004 within the framework of the Artificial Intelligence National Laboratory of Hungary.
5 Proofs
5.1 Proof of Theorem 2.1
Proof.
It is known that
$L(D)-L^{*}$ $=\int{\mathbb{I}}_{\{D(x)\neq D^{*}(x)\}}|m(x)|\mu(dx),$ (17)
cf. Theorem 2.2 in Devroye, Györfi and Lugosi [6]. For notational simplicity let $h=h_{n}$. For $x\in A_{h,j}$, we set
$m_{n}(x)=\frac{\nu_{n}(A_{h,j})}{\mu(A_{h,j})}.$
Because of (17), we have that
$L(D_{n})-L^{*}$ $=\int{\mathbb{I}}_{\{\mbox{sign }(m_{n}(x))\neq\mbox{sign }(m(x))\}}|m(x)|\mu(dx)$
$\leq\int{\mathbb{I}}_{\{|m_{n}(x)-m(x)|\geq|m(x)|\}}|m(x)|\mu(dx)$
$\leq I_{n}+J_{n},$
where
$I_{n}$ $=\int{\mathbb{I}}_{\{|{\mathbb{E}}\{m_{n}(x)\}-m(x)|\geq|m(x)|/2\}}|m(x)|\mu(dx)$
and
$J_{n}$ $=\int{\mathbb{I}}_{\{|m_{n}(x)-{\mathbb{E}}\{m_{n}(x)\}|\geq|m(x)|/2\}}|m(x)|\mu(dx).$
As the approximation error $I_{n}$,
$I_{n}$ $=\sum_{j}\int_{A_{h,j}}{\mathbb{I}}_{\{|\nu(A_{h,j})-\mu(A_{h,j})m(x)|\geq\mu(A_{h,j})|m(x)|/2\}}|m(x)|\mu(dx),$
where $\nu(A_{h,j})={\mathbb{E}}[Y{\mathbb{I}}_{\{X\in A_{h,j}\}}]$. The Lipschitz condition implies that
$|\nu(A_{h,j})-\mu(A_{h,j})m(x)|$ $\leq\left|\int_{A_{h,j}}m(z)\mu(dz)-\mu(A_{h,j})m(x)\right|$
$\leq\int_{A_{h,j}}\left|m(z)-m(x)\right|\mu(dz)$
$\leq C\sqrt{d}h\mu(A_{h,j}).$
This together with the margin condition yields that
$I_{n}$ $\leq\sum_{j}\int_{A_{h,j}}{\mathbb{I}}_{\{C\sqrt{d}h\geq|m(x)|/2\}}|m(x)|\mu(dx)\leq c^{*}(2C\sqrt{d}h)^{1+\gamma},$ (18)
and thus the bound on the approximation error in (11).
Concerning the estimation error $J_{n}$, we have that
${\mathbb{E}}\{J_{n}\}$ $=\sum_{A\in{\cal P}_{h}}\int_{A}{\mathbb{P}}\{|m_{n}(x)-{\mathbb{E}}\{m_{n}(x)\}|\geq|m(x)|/2\}|m(x)|\mu(dx)$
$=\sum_{A\in{\cal P}_{h}}\int_{A}{\mathbb{P}}\{|\nu_{n}(A)-\nu(A)|\geq\mu(A)|m(x)|/2\}|m(x)|\mu(dx).$
We have that
$\sum_{A\in{\cal P}_{h}}\int_{A}{\mathbb{P}}\Big{(}|\nu_{n}(A)-\nu(A)|\geq\mu(A)|m(x)|/2\Big{)}|m(x)|\mu(dx)$
$\approx 2\sum_{A\in{\cal P}_{h}}\int_{A}\Phi\left(-\sqrt{n}\frac{\mu(A)|m(x)|/2}{\sqrt{{\mathbb{V}ar}(Y{\mathbb{I}}_{\{X\in A\}}})}\,\right)|m(x)|\mu(dx),$ (19)
where in the last step we refer to CLT. (At the end of this subsection we show that the error term for this CLT approximation is of order $O(1/(nh_{n}^{d_{a}}))$.)
Because of ${\mathbb{V}ar}(Y{\mathbb{I}}_{\{X\in A\}})\leq{\mathbb{E}}[{\mathbb{I}}_{\{X\in A\}}]=\mu(A)$ and up to the error term just mentioned, this implies
${\mathbb{E}}\{J_{n}\}$ $\leq 2\sum_{A\in{\cal P}_{h}}\int_{A}\Phi\left(-\sqrt{n}\sqrt{\mu(A)}|m(x)|/2\,\right)|m(x)|\mu(dx).$ (20)
Next, we use the inequalities
$\Phi(-t)\leq e^{-t^{2}/2}\frac{1}{\sqrt{2\pi}}\frac{1}{t},$
($t>0$, cf. p. 179 in Feller [10]). This together with (20) implies
${\mathbb{E}}\{J_{n}\}$ $\leq 2\sum_{A\in{\cal P}_{h}}\int_{A}\exp\left(-\frac{n}{8}\mu(A)m(x)^{2}\right)\frac{|m(x)|}{\sqrt{n}|m(x)|\sqrt{\mu(A)}/2}\mu(dx)$
$\leq\frac{4}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}\int_{A}\exp\left(-\frac{n}{8}\mu(A)m(x)^{2}\right)\frac{1}{\sqrt{\mu(A)}}\mu(dx).$
From the decomposition (7) one gets that
${\mathbb{E}}\{J_{n}\}$ $\leq\frac{4}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{\mu(A)>h^{d_{a}}\}}\int_{A}\exp\left(-\frac{n}{8}m(x)^{2}\mu(A)\right)\mu(dx)\frac{1}{\sqrt{\mu(A)}}$
$+\frac{4}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu(A)>0\}}\int_{A}\exp\left(-\frac{n}{8}m(x)^{2}\mu(A)\right)\mu_{a}(dx)\frac{1}{\sqrt{\mu(A)}}$
$+\frac{4}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu_{s}(A)>0\}}\sqrt{\mu_{s}(A)}.$
The discrete part is easy. The set $S_{s}$ is finite, therefore
$\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu_{s}(A)>0\}}$ $\leq\sum_{x\in S_{s}}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu_{s}(\{x\})>0\}}=0$ (21)
for $h$ small enough. (We note that the CLT approximation is not needed for the discrete part.)
Additionally, one has that
$\frac{4}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{\mu(A)>h^{d_{a}}\}}\int_{A}\exp\left(-\frac{n}{8}m(x)^{2}\mu(A)\right)\mu(dx)\frac{1}{\sqrt{\mu(A)}}$
$\leq\frac{4}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{\mu(A)>h^{d_{a}}\}}\int_{A}\exp\left(-\frac{nh^{d_{a}}}{8}m(x)^{2}\right)\mu(dx)\frac{1}{\sqrt{h^{d_{a}}}}$
$\leq\frac{4}{\sqrt{nh^{d_{a}}}}\int\exp\left(-\frac{nh^{d_{a}}}{8}m(x)^{2}\right)\mu(dx).$
For $a=nh^{d_{a}}/8$, partial integration together with the margin condition implies
$\frac{4}{\sqrt{nh^{d_{a}}}}\int\exp\left(-nh^{d_{a}}m(x)^{2}/8\right)\mu(dx)$ $=\frac{4}{\sqrt{nh^{d_{a}}}}\int_{0}^{\infty}e^{-as^{2}}G^{*}(ds)$
$=\frac{4}{\sqrt{nh^{d_{a}}}}2a\int_{0}^{\infty}se^{-as^{2}}G^{*}(s)ds$
$\leq\frac{4}{\sqrt{nh^{d_{a}}}}2c^{*}a\int_{0}^{\infty}e^{-as^{2}}s^{1+\gamma}ds$
$=\frac{4}{\sqrt{nh^{d_{a}}}}2c^{*}a^{-\gamma/2}\int_{0}^{\infty}e^{-u^{2}}u^{1+\gamma}du$
$=O\left(\frac{1}{(nh^{d_{a}})^{(\gamma+1)/2}}\right),$ (22)
where recall that $G^{*}(s)=\int{\mathbb{I}}_{\{0<|m(x)|\leq t\}}\mu(dx)$. (Note that the CLT approximation with an error term $O(1/(nh_{n}^{d_{a}}))$ is justified by (26) below.)
For notation (8) and for $a=nh^{d_{a}}/8$, the combined margin and density assumption implies that for $0<h\leq h_{1}^{*}$, we have
$\frac{4}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu(A)>0\}}\int_{A}\exp\left(-\frac{n}{8}m(x)^{2}\mu(A)\right)\mu_{a}(dx)\frac{1}{\sqrt{\mu(A)}}$
$\leq\frac{4}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}\int_{A}\exp\left(-\frac{n}{8}m(x)^{2}\mu_{a}(A)\right)\mu_{a}(dx)\frac{1}{\sqrt{\mu_{a}(A)}}$
$=\frac{4}{\sqrt{n}}\int_{S_{a}}\exp\left(-nh^{d_{a}}m(x)^{2}f_{h}(x)/8\right)\frac{1}{\sqrt{h^{d_{a}}f_{h}(x)}}\mu_{a}(dx)$
$=\frac{4}{\sqrt{nh^{d_{a}}}}\int_{0}^{\infty}e^{-as^{2}}G_{h}(ds)$
$=\frac{4}{\sqrt{nh^{d_{a}}}}2a\int_{0}^{\infty}se^{-as^{2}}G_{h}(s)ds$
$\leq\frac{4}{\sqrt{nh^{d_{a}}}}2c_{1}a\int_{0}^{\infty}e^{-as^{2}}s^{1+\gamma_{1}}ds$
$=\frac{4}{\sqrt{nh^{d_{a}}}}2c_{1}a^{-\gamma_{1}/2}\int_{0}^{\infty}e^{-u^{2}}u^{1+\gamma_{1}}du$
$=O\left(\frac{1}{(nh_{n}^{d_{a}})^{(\gamma_{1}+1)/2}}\right),$ (23)
(We note that the CLT approximation with an error term $O(1/(nh_{n}^{d_{a}}))$ is justified by (25) below.) (21), (5.1) and (23) together with (25) and (26) below yield the bound on the estimation error in (11). ∎
The CLT approximation error.
For the CLT approximation in (5.1), we need an upper bound. Put $Z_{i,A}=Y_{i}{\mathbb{I}}_{\{X_{i}\in A\}}$. We use the Berry-Esséen theorem and the normal approximation to upper bound this probability. Recall that because of the nonuniform Berry-Esséen theorem there exists a universal constant $0.4097<c<0.4785$ such that for all $a\in{\mathbb{R}}$ we have
$\left|{\mathbb{P}}\Big{(}\,\frac{\sqrt{n}}{\sigma_{A}n}\sum_{i=1}^{n}Z_{i,A}-\frac{\sqrt{n}}{\sigma_{A}}{\mathbb{E}}Z_{i,A}<a\,\Big{)}-\Phi(a)\right|\leq\frac{c\varrho_{A}}{(1+|a|^{3})\sigma_{A}^{3}\sqrt{n}},$
where $\varrho_{A}={\mathbb{E}}[|Z_{i,A}-{\mathbb{E}}Z_{i,A}|^{3}]$ and $\sigma^{2}_{A}={\mathbb{V}ar}(Z_{i,A})$, see Esséen [9] and Tyurin [18]. Thus,
${\mathbb{P}}\bigg{(}\,\frac{\sqrt{n}}{\sigma_{A}n}\sum_{i=1}^{n}Z_{i,A}-\frac{\sqrt{n}}{\sigma_{A}}{\mathbb{E}}Z_{i,A}<a\,\bigg{)}$ $\leq\Phi(a)+\frac{c\varrho_{A}}{(1+|a|^{3})\sigma_{A}^{3}\sqrt{n}}.$
It implies that
${\mathbb{P}}\left(\,\nu_{n}(A)-\nu(A)\geq\mu(A)|m(x)|/2\,\right)$
$={\mathbb{P}}\left(\,\frac{\sqrt{n}}{\sigma_{A}}\nu_{n}(A)-\frac{\sqrt{n}}{\sigma_{A}}\nu(A)\geq\frac{\sqrt{n}}{\sigma_{A}}\mu(A)|m(x)|/2\,\right)$
$\leq\Phi\left(\,-\sqrt{n}\frac{\mu(A)|m(x)|/2}{\sigma_{A}}\,\right)+\frac{c\varrho_{A}}{(1+|\sqrt{n}\frac{\mu(A)|m(x)|/2}{\sigma_{A}}|^{3})\sigma_{A}^{3}\sqrt{n}}$
Therefore, the error term in the approximation (5.1) is equal to
$\Delta$ $:=2\sum_{A\in{\cal P}_{h}}\int_{A}\frac{c\varrho_{A}}{(1+|\sqrt{n}\frac{\mu(A)|m(x)|/2}{\sigma_{A}}|^{3})\sigma_{A}^{3}\sqrt{n}}|m(x)|\mu(dx)$
$\leq 4\sum_{A\in{\cal P}_{h}}\int_{A}\frac{c\varrho_{A}}{\sigma_{A}^{2}n}\frac{\sqrt{n}\mu(A)|m(x)|/(2\sigma_{A})}{(1+|\sqrt{n}\mu(A)|m(x)|/(2\sigma_{A})|^{3})}\mu(dx)\frac{1}{\mu(A)}$
$\leq 4\max_{z}\frac{z}{1+z^{3}}\sum_{A\in{\cal P}_{h}}\frac{c\varrho_{A}}{\sigma_{A}^{2}n}.$ (24)
A simple consideration yields
$\varrho_{A}$ $=8{\mathbb{E}}\left[\left(\frac{|Z_{i,A}-{\mathbb{E}}Z_{i,A}|}{2}\right)^{3}\right]\leq 8{\mathbb{E}}\left[\left(\frac{|Z_{i,A}-{\mathbb{E}}Z_{i,A}|}{2}\right)^{2}\right]=2\sigma_{A}^{2}.$
Therefore, by the boundedness of $X$,
$\Delta$ $\leq O\left(\frac{1}{n}\right)\sum_{A\in{\cal P}_{h}}1=O\left(\frac{1}{nh_{n}^{d}}\right).$
If $\Delta$ is replaced by $\Delta^{\prime}$ via replacing $\mu(dx)$ by $\mu_{a}(dx)$, then
$\Delta^{\prime}$ $=O\left(\frac{1}{n}\right)\sum_{A\in{\cal P}_{h}}\frac{1}{\mu(A)}\mu_{a}(A)\leq O\left(\frac{1}{n}\right)\sum_{A\in{\cal P}_{h}}1=O\left(\frac{1}{nh_{n}^{d_{a}}}\right).$ (25)
If $\Delta$ is modified by $\Delta^{\prime\prime}$ via inserting the factor ${\mathbb{I}}_{\{\mu(A)>h^{d_{a}}\}}\leq\mu(A)/h^{d_{a}}$ into the summands, then
$\Delta^{\prime\prime}$ $=O\left(\frac{1}{n}\right)\sum_{A\in{\cal P}_{h}}\frac{1}{h^{d_{a}}}\mu(A)=O\left(\frac{1}{nh_{n}^{d_{a}}}\right).$ (26)
5.2 Proof of Lemma 2.2
Proof.
One has that
$G_{h}(t)$ $=\int_{S_{a}}{\mathbb{I}}_{\{0<{\sqrt{f_{h}(x)}|m(x)|\leq t\}}}\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx)$
$=\int_{B^{*}_{\epsilon}}{\mathbb{I}}_{\{0<{\sqrt{f_{h}(x)}|m(x)|\leq t\}}}\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx)$
$+\int_{S_{a}\setminus B^{*}_{\epsilon}}{\mathbb{I}}_{\{0<{\sqrt{f_{h}(x)}|m(x)|\leq t\}}}\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx)$
$\leq\int_{B^{*}_{\epsilon}}{\mathbb{I}}_{\{0<{\sqrt{f_{\epsilon,min}}|m(x)|\leq t\}}}\frac{1}{\sqrt{f_{\epsilon,min}}}\mu_{a}(dx)$
$+\int_{S_{a}\setminus B^{*}_{\epsilon}}{\mathbb{I}}_{\{0<{\sqrt{f_{h}(x)}\epsilon\leq t\}}}\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx).$
By the definition of the margin condition,
$\int_{B^{*}_{\epsilon}}{\mathbb{I}}_{\{0<{\sqrt{f_{\epsilon,min}}|m(x)|\leq t\}}}\frac{1}{\sqrt{f_{\epsilon,min}}}\mu_{a}(dx)$ $\leq\int{\mathbb{I}}_{\{0<|m(x)|\leq t/\sqrt{f_{\epsilon,min}}\}}\mu(dx)/\sqrt{f_{\epsilon,min}}$
$=G^{*}\left(t/\sqrt{f_{\epsilon,min}}\right)/\sqrt{f_{\epsilon,min}}$
$\leq c^{*}\left(t/\sqrt{f_{\epsilon,min}}\right)^{\gamma}/\sqrt{f_{\epsilon,min}}.$
For the notation
$H(s)$ $:=\mu_{a}(\{x:0<f_{h}(x)\leq s\}),$
one gets that
$H(s)$ $=\sum_{j}\mu_{a}(\{x:0<f_{h}(x)\leq s,x\in A_{h,j}\})$
$=\sum_{j}\mu_{a}(\{x:0<\mu_{a}(A_{h,j})/h^{d_{a}}\leq s,x\in A_{h,j}\})$
$=\sum_{j}{\mathbb{I}}_{\{0<\mu_{a}(A_{h,j})/h^{d_{a}}\leq s\}}\mu_{a}(A_{h,j})$
$\leq\sum_{j}{\mathbb{I}}_{\{0<\mu_{a}(A_{h,j})\}}h^{d_{a}}\cdot s$
$\leq const\cdot s.$
Therefore,
$\int_{S_{a}\setminus B^{*}_{\epsilon}}{\mathbb{I}}_{\{0<{\sqrt{f_{h}(x)}\epsilon\leq t\}}}\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx)$ $\leq\int_{S_{a}}{\mathbb{I}}_{\{0<{\sqrt{f_{h}(x)}\leq t/\epsilon\}}}\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx)$
$=\int_{0}^{(t/\epsilon)^{2}}\frac{1}{\sqrt{s}}H(ds)$
$\leq const\cdot t/\epsilon,$
and the lemma is proved. ∎
5.3 Proof of Theorem 2.3
Let
$P_{n,k}(x)=\frac{\nu_{n,k}(A_{h,j})}{\mu(A_{h,j})},\quad\mbox{ if }x\in A_{h,j}.$
The main ingredient of the proofs is the slight extension of Lemma 8 in [13]:
Lemma 5.1.
Let $g_{n}$ be a plug-in rule with any estimates $P_{n,j}$ of $P_{j}$. For the notation
$\Delta_{l}(x)$ $=P_{(1)}(x)-P_{(l)}(x),$
we have
${\mathbb{E}}\{L(g_{n})\}-L^{*}$ $\leq\sum_{j=1}^{M}\sum_{l=2}^{M}J_{n,j,l}$
where
$J_{n,j,l}=\int\Delta_{l}(x){\mathbb{P}}\{|P_{n,j}(x)-P_{j}(x)|\geq\Delta_{l}(x)/2\}\mu(dx).$ (27)
Proof.
As in the proof of Lemma 8 in Györfi and Weiss [13],
${\mathbb{E}}\{L(g_{n})\}-L^{*}$
$=\int{\mathbb{E}}\{(P_{g^{*}(x)}(x)-P_{g_{n}(x)}(x)){\mathbb{I}}_{\{g^{*}(x)\neq g_{n}(x)\}}\}\mu(dx)$
$=\int{\mathbb{E}}\{(P_{g^{*}(x)}(x)-P_{g_{n}(x)}(x)){\mathbb{I}}_{\{P_{g^{*}(x)}(x)>P_{g_{n}(x)}(x)\}}{\mathbb{I}}_{\{P_{n,g_{n}(x)}(x)\geq P_{n,g^{*}(x)}(x)\}}\}\mu(dx).$
If $g^{*}(x)=j$ and $g_{n}(x)=l$, then
$\{P_{n,g_{n}(x)}(x)-P_{n,g^{*}(x)}(x)\geq 0\}$
$=\{P_{n,l}(x)-P_{l}(x)+P_{l}(x)-P_{g^{*}(x)}(x)+P_{j}(x)-P_{n,j}(x)\geq 0\}$
$\subset\left\{|P_{n,j}(x)-P_{j}(x)|\geq(P_{g^{*}(x)}(x)-P_{g_{n}(x)}(x))/2\right\}$
$\cup\left\{|P_{n,l}(x)-P_{l}(x)|\geq(P_{g^{*}(x)}(x)-P_{g_{n}(x)}(x))/2\right\}$
Therefore,
${\mathbb{E}}\{L(g_{n})\}-L^{*}$
$\leq\sum_{j=1}^{M}\int{\mathbb{E}}\{(P_{g^{*}(x)}(x)-P_{g_{n}(x)}(x)){\mathbb{I}}_{\{|P_{n,j}(x)-P_{j}(x)|\geq(P_{g^{*}(x)}(x)-P_{g_{n}(x)}(x))/2\}}\}\mu(dx)$
$\leq\sum_{j=1}^{M}\sum_{l=2}^{M}\int(P_{(1)}(x)-P_{(l)}(x)){\mathbb{E}}\{{\mathbb{I}}_{\{|P_{n,j}(x)-P_{j}(x)|\geq(P_{(1)}(x)-P_{(l)}(x))/2\}}\}\mu(dx)$
$=\sum_{j=1}^{M}\sum_{l=2}^{M}\int\Delta_{l}(x){\mathbb{P}}\{|P_{n,j}(x)-P_{j}(x)|\geq\Delta_{l}(x)/2\}\mu(dx).$
Proof.
As above let us use the simplified notation $h=h_{n}$. We bound (27) by
$J_{n,k,l}\leq J_{n,k,l}^{(1)}+J_{n,k,l}^{(2)}\,$,
where
$J_{n,k,l}^{(1)}$ $=\int\Delta_{l}(x){\mathbb{P}}\{|P_{n,k}(x)-{\mathbb{E}}\{P_{n,k}(x)\}|\geq\Delta_{l}(x)/4\}\mu(dx),$
and
$J_{n,k,l}^{(2)}$ $=\int\Delta_{l}(x){\mathbb{I}}_{\{|{\mathbb{E}}\{P_{n,k}(x)\}-P_{k}(x)|\geq\Delta_{l}(x)/4\}}\mu(dx).$
Concerning the estimation error $J_{n,k,l}^{(1)}$, as in the proof of Theorem 2.1 we apply the CLT with the Berry-Esséen bound and then decomposition (7) and also (21). For $x\in A_{h,j}$, we have that
${\mathbb{P}}\{|P_{n,k}(x)-{\mathbb{E}}\{P_{n,k}(x)\}|\geq\Delta_{l}(x)/4\}$
$={\mathbb{P}}\{|\nu_{n,k}(A_{h,j})-{\mathbb{E}}\{\nu_{n,k}(A_{h,j})\}|\geq\mu(A_{h,j})\Delta_{l}(x)/4\}$
$\approx 2\Phi\left(-\sqrt{n}\frac{\mu(A_{h,j})|\Delta_{l}(x)|/4}{\sqrt{{\mathbb{V}ar}(Y{\mathbb{I}}_{\{X\in A_{h,j}\}}})}\,\right),$
and henceforth, up to a term $O(1/(nh_{n}^{d_{a}}))$,
$\sum_{A\in{\cal P}_{h}}\int_{A}{\mathbb{P}}\Big{(}|\nu_{n,k}(A)-{\mathbb{E}}[\nu_{n,k}(A)]|\geq\mu(A)|\Delta_{l}(x)|/4\Big{)}|\Delta_{l}(x)|\mu(dx)$
$\approx 2\sum_{A\in{\cal P}_{h}}\int_{A}\Phi\left(-\sqrt{n}\frac{\mu(A)|\Delta_{l}(x)|/4}{\sqrt{{\mathbb{V}ar}(Y{\mathbb{I}}_{\{X\in A\}}})}\,\right)|\Delta_{l}(x)|\mu(dx)$
$\leq\frac{8}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}\int_{A}\exp\left(-\frac{n}{32}\mu(A)\Delta_{l}(x)^{2}\right)\frac{1}{\sqrt{\mu(A)}}\mu(dx)$
$\leq\frac{8}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{\mu(A)>h^{d_{a}}\}}\int_{A}\exp\left(-\frac{n}{32}\Delta_{l}(x)^{2}\mu(A)\right)\mu(dx)\frac{1}{\sqrt{\mu(A)}}$
$\quad+\frac{8}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu_{a}(A)>0\}}\int_{A}\exp\left(-\frac{n}{32}\Delta_{l}(x)^{2}\mu_{a}(A)\right)\mu_{a}(dx)\frac{1}{\sqrt{\mu(A)}}$
$\quad+\frac{8}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu_{s}(A)>0\}}\sqrt{\mu_{s}(A)},$
which is smaller than
$\frac{8}{\sqrt{nh^{d_{a}}}}\int\exp\left(-\frac{nh^{d_{a}}}{32}\Delta_{l}(x)^{2}\right)\mu(dx)$
$+\frac{8}{\sqrt{nh^{d_{a}}}}\int_{S_{a}}\exp\left(-\frac{nh^{d_{a}}}{32}\Delta_{l}(x)^{2}f_{h}(x)\right)\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx),$
for $h$ small enough. If $l\neq g^{*}(x)$, then $\Delta_{l}(x)\geq P_{(1)}(x)-P_{(2)}(x)=:\Delta(x)$ otherwise $\Delta_{l}(x)=0$, therefore up to a Berry-Esséen error term we have
$J_{n,k,l}^{(1)}$ $\leq\frac{8}{\sqrt{nh^{d_{a}}}}\int\exp\left(-\frac{nh^{d_{a}}}{32}\Delta(x)^{2}\right)\mu(dx)$
$\quad+\frac{8}{\sqrt{nh^{d_{a}}}}\int_{S_{a}}\exp\left(-\frac{nh^{d_{a}}}{32}\Delta(x)^{2}f_{h}(x)\right)\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx).$
and also
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(1)}$ $\leq\frac{8M^{2}}{\sqrt{nh^{d_{a}}}}\int\exp\left(-\frac{nh^{d_{a}}}{32}\Delta(x)^{2}\right)\mu(dx)$
$\quad+\frac{8M^{2}}{\sqrt{nh^{d_{a}}}}\int_{S_{a}}\exp\left(-\frac{nh^{d_{a}}}{32}\Delta(x)^{2}f_{h}(x)\right)\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx).$ (28)
Similarly to (5.1), the margin condition implies
$\frac{8M^{2}}{\sqrt{nh^{d_{a}}}}\int\exp\left(-\frac{nh^{d_{a}}}{32}(P_{(1)}(x)-P_{(2)}(x))^{2}\right)\mu(dx)$
$=\frac{8M^{2}}{\sqrt{nh^{d_{a}}}}\int_{0}^{\infty}e^{-as^{2}}G^{*}(ds)$
$=O\left((nh_{n}^{d_{a}})^{-(\gamma+1)/2}\right),$ (29)
where $a=(nh^{d_{a}})/32$. As (23), the combined margin and density condition yields that
$\frac{8M^{2}}{\sqrt{nh^{d_{a}}}}\int_{S_{a}}\exp\left(-\frac{nh^{d_{a}}}{32}(P_{(1)}(x)-P_{(2)}(x))^{2}f_{h}(x)\right)\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx)$
$=\frac{8M^{2}}{\sqrt{nh^{d_{a}}}}\int_{0}^{\infty}e^{-as^{2}}G_{h}(ds)$
$=O\left((nh_{n}^{d_{a}})^{-(\gamma_{1}+1)/2}\right).$ (30)
Thus,
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(1)}$ $=O\left((nh_{n}^{d_{a}})^{-(\gamma_{1}+1)/2}\right)+O\left((nh_{n}^{d_{a}})^{-(\gamma+1)/2}\right)+O\bigg{(}\frac{1}{nh_{n}^{d_{a}}}\bigg{)}.$ (31)
Concerning the approximation error $J_{n,j,l}^{(2)}$ (compare the proof of (18)), we have
$J_{n,k,l}^{(2)}$ $=\sum_{j}\int_{A_{h,j}}\Delta_{l}(x){\mathbb{I}}_{\{|{\mathbb{E}}\{P_{n,k}(x)\}-P_{k}(x)|\geq\Delta_{l}(x)/4\}}\mu(dx)$
$=\sum_{j}\int_{A_{h,j}}\Delta_{l}(x){\mathbb{I}}_{\{|\nu_{k}(A_{h,j})-\mu(A_{h,j})P_{k}(x)|\geq\mu(A_{h,j})\Delta_{l}(x)/4\}}|m(x)|\mu(dx).$
The Lipschitz condition implies that
$|\nu_{k}(A_{h,j})-\mu(A_{h,j})P_{k}(x)|$ $\leq\left|\int_{A_{h,j}}P_{k}(z)\mu(dz)-\mu(A_{h,j})P_{k}(x)\right|$
$\leq\int_{A_{h,j}}\left|P_{k}(z)-P_{k}(x)\right|\mu(dz)$
$\leq C\sqrt{d}h\mu(A_{h,j}),$
where $\nu_{k}(A_{h,j})=\int_{A_{h,j}}P_{k}(z)\mu(dz)$. This together with the margin condition yields that
$J_{n,k,l}^{(2)}$ $\leq\sum_{j}\int_{A_{h,j}}{\mathbb{I}}_{\{C\sqrt{d}h\geq\Delta_{l}(x)/4\}}\Delta_{l}(x)\mu(dx)\leq c^{*}(4C\sqrt{d}h)^{1+\gamma},$ (32)
and so
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(2)}$ $=M^{2}O\left(h_{n}^{1+\gamma}\right).$ (33)
5.4 Proof of Theorem 3.1
Proof.
Let $h=h_{n}$ be as above. For $x\in A_{h,j}$, we set
${\widehat{m}}_{n}(x)=\frac{{\widetilde{\nu}}_{n}(A_{h,j})}{\mu(A_{h,j})},$
and
${{\widehat{m}}^{\prime}}_{n}(x)=\frac{\frac{\sigma_{Z}}{n}\sum_{i=1}^{n}\epsilon_{i,j}}{\mu(A_{h,j})}.$
Then,
${\widehat{m}}_{n}={{\widehat{m}}^{\prime}}_{n}+m_{n},$
and
$\widetilde{D}_{n}(x)=\mbox{sign }({\widehat{m}}_{n}(x)).$
Because of (17), we have that
$L(\widetilde{D}_{n})-L^{*}$ $=\int{\mathbb{I}}_{\{\mbox{sign }({\widehat{m}}_{n}(x))\neq\mbox{sign }(m(x))\}}|m(x)|\mu(dx)$
$\leq\int{\mathbb{I}}_{\{|{\widehat{m}}_{n}(x)-m(x)|\geq|m(x)|\}}|m(x)|\mu(dx)$
$\leq I_{n}+J_{n}+K_{n},$
where
$I_{n}$ $=\int{\mathbb{I}}_{\{|{\mathbb{E}}\{m_{n}(x)\}-m(x)|\geq|m(x)|/3\}}|m(x)|\mu(dx),$
and
$J_{n}$ $=\int{\mathbb{I}}_{\{|m_{n}(x)-{\mathbb{E}}\{m_{n}(x)\}|\geq|m(x)|/3\}}|m(x)|\mu(dx),$
and
$K_{n}$ $=\int{\mathbb{I}}_{\{|{{\widehat{m}}^{\prime}}_{n}(x)|\geq|m(x)|/3\}}|m(x)|\mu(dx).$
As in the proof of Theorem 2.1 one gets that
$I_{n}+{\mathbb{E}}\{J_{n}\}$ $=O\left(h_{n}^{1+\gamma}\right)+O\left(\bigg{(}\frac{1}{nh_{n}^{d_{a}}}\bigg{)}^{(\min\{1,\gamma,\gamma_{1})\}+1)/2}\right).$
For the term $K_{n}$,
${\mathbb{E}}\{K_{n}\}$ $=\int{\mathbb{P}}\left\{|{{\widehat{m}}^{\prime}}_{n}(x)|\geq|m(x)|/3\right\}|m(x)|\mu(dx)$
$=\sum_{j}\int_{A_{h,j}}{\mathbb{P}}\left\{\left|\frac{\sigma_{Z}}{n}\sum_{i=1}^{n}\epsilon_{i,j}\right|\geq\mu(A_{h,j})|m(x)|/3\right\}|m(x)|\mu(dx).$
The Berry-Esséen theorem as above yields
${\mathbb{P}}\left\{\left|\frac{\sigma_{Z}}{n}\sum_{i=1}^{n}\epsilon_{i,j}\right|\geq\mu(A_{h,j})|m(x)|/3\right\}$
$=2\cdot{\mathbb{P}}\left\{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\epsilon_{i,j}\geq\frac{\sqrt{n}\mu(A_{h,j})|m(x)|}{3\sigma_{Z}}\right\}$
$\leq 2\Phi\bigg{(}-\frac{\sqrt{n}\mu(A_{h,j})|m(x)|}{3\sigma_{Z}}\bigg{)}+\frac{2c\varrho_{\varepsilon}}{\sigma_{\varepsilon}^{3}\Big{(}1+\Big{|}\frac{\sqrt{n}\mu(A_{h,j})|m(x)|}{3\sigma_{Z}}\Big{|}^{3}\Big{)}\sqrt{n}}.$ (34)
Similarly as in (5.1) the error of CLT is essentially dominated by the first term in (5.4), since
$\Delta$ $=2\sum_{j}\int_{A_{h,j}}\frac{c\varrho_{\varepsilon}}{\sigma_{\varepsilon}^{3}\Big{(}1+\Big{|}\frac{\sqrt{n}\mu(A_{h,j})|m(x)|}{3\sigma_{Z}}\Big{|}^{3}\Big{)}\sqrt{n}}|m(x)|\mu(dx)$
$\leq 6\max_{z}\frac{z}{1+z^{3}}\sum_{j}\frac{c\varrho_{\varepsilon}\sigma_{Z}}{\sigma_{\varepsilon}^{3}n}=\frac{\tilde{c}\sigma_{Z}}{nh^{d}}$
by the boundedness of $X$. Using decomposition (7), as in the proof of Theorem 2.1 with (21), (5.1) and (23) one obtains
$2\sum_{j}\int_{A_{h,j}}\Phi\bigg{(}-\frac{\sqrt{n}\mu(A_{h,j})|m(x)|}{3\sigma_{Z}}\bigg{)}|m(x)|\mu(dx)$
$\leq 2\sum_{j}\int_{A_{h,j}}\frac{3\sigma_{Z}}{\sqrt{n}\mu(A_{h,j})|m(x)|}\exp\bigg{(}-\frac{n\mu(A_{h,j})^{2}m(x)^{2}}{18\sigma_{Z}^{2}}\bigg{)}|m(x)|\mu(dx).$
$\leq 2\sum_{j}{\mathbb{I}}_{\{h^{d_{a}}<\mu(A_{h,j})\}}\int_{A_{h,j}}\frac{3}{\sqrt{n\mu(A_{h,j})^{2}/\sigma_{Z}^{2}}}\exp\left(-\frac{n\mu(A_{h,j})^{2}m(x)^{2}}{18\sigma_{Z}^{2}}\right)\mu(dx)$
$\quad+2\sum_{j}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu(A_{h,j})>0\}}\int_{A_{h,j}}\frac{3}{f_{h}(x)\sqrt{nh^{2d_{a}}/\sigma_{Z}^{2}}}\exp\left(-\frac{nf_{h}(x)^{2}h^{2d_{a}}m(x)^{2}}{18\sigma_{Z}^{2}}\right)\mu_{a}(dx)$
$\leq 2\int\frac{3}{\sqrt{nh^{2d_{a}}/\sigma_{Z}^{2}}}\exp\left(-\frac{nh^{2d_{a}}m(x)^{2}}{18\sigma_{Z}^{2}}\right)\mu(dx)$
$\quad+2\int\frac{3}{f_{h}(x)\sqrt{nh^{2d_{a}}/\sigma_{Z}^{2}}}\exp\left(-\frac{nf_{h}(x)^{2}h^{2d_{a}}m(x)^{2}}{18\sigma_{Z}^{2}}\right)\mu_{a}(dx)$
for $h$ small enough. By the margin condition
$2\int\frac{3}{\sqrt{nh^{2d_{a}}/\sigma_{Z}^{2}}}\exp\left(-\frac{nh^{2d_{a}}m(x)^{2}}{18\sigma_{Z}^{2}}\right)\mu(dx)$ $=O\left(\left(\frac{\sigma_{Z}^{2}}{nh_{n}^{2d_{a}}}\right)^{(\gamma+1)/2}\right).$
and by the modified combined margin and density condition
$\int\frac{3}{f_{h}(x)\sqrt{nh^{2d_{a}}/\sigma_{Z}^{2}}}\exp\left(-\frac{nf_{h}(x)^{2}h^{2d_{a}}m(x)^{2}}{18\sigma_{Z}^{2}}\right)\mu_{a}(dx)$
$=O\left(\left(\frac{\sigma_{Z}^{2}}{nh_{n}^{2d_{a}}}\right)^{(\gamma_{2}+1)/2}\right).$
In conclusion, we have
${\mathbb{E}}(L(\widetilde{g}_{n}))-L^{*}=O\left(h^{1+\gamma}\right)+O\left(\bigg{(}\frac{1}{nh_{n}^{d_{a}}}\bigg{)}^{(\min\{1,\gamma,\gamma_{1})\}+1)/2}\right)$
$\quad+O\left(\left(\frac{\sigma_{Z}^{2}}{nh^{2d_{a}}}\right)^{(\gamma+1)/2}\right)+O\left(\left(\frac{\sigma_{Z}^{2}}{nh^{2d_{a}}}\right)^{(\gamma_{2}+1)/2}\right)+O\left(\frac{\sigma_{Z}}{nh^{d}}\right).$
5.5 Proof of Theorem 3.2
Proof.
Again, we bound (27) by
$J_{n,k,l}\leq J_{n,k,l}^{(1)}+J_{n,k,l}^{(2)}+J_{n,k,l}^{(3)},$
where
$J_{n,k,l}^{(1)}$ $=\int\Delta_{l}(x){\mathbb{P}}\{|P_{n,k}(x)-{\mathbb{E}}\{P_{n,k}(x)\}|\geq\Delta_{l}(x)/6\}\mu(dx),$
and
$J_{n,k,l}^{(2)}$ $=\int\Delta_{l}(x){\mathbb{I}}_{\{|{\mathbb{E}}\{P_{n,k}(x)\}-P_{k}(x)|\geq\Delta_{l}(x)/6\}}\mu(dx),$
and
$J_{n,k,l}^{(3)}$ $=\int\Delta_{l}(x){\mathbb{P}}\{|Q_{n,k}(x)|\geq\Delta_{l}(x)/6\}\mu(dx),$
with
$Q_{n,k}(x)=\frac{\frac{\sigma_{Z}}{n}\sum_{i=1}^{n}\epsilon_{i,j,k}}{\mu(A_{h,j})},\quad\mbox{ if }x\in A_{h,j}.$
As in (31) and (33), up to a Berry-Esséen bound
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(1)}$ $=O\left((nh_{n}^{d_{a}})^{-(\gamma_{1}+1)/2}\right)+O\left((nh_{n}^{d_{a}})^{-(\gamma+1)/2}\right),$
and
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(2)}$ $=O\left(h_{n}^{1+\gamma}\right).$
Again, the CLT yields that
$\Delta_{l}(x){\mathbb{P}}\left\{\left|\frac{\sigma_{Z}}{n}\sum_{i=1}^{n}\epsilon_{i,j,k}\right|\geq\mu(A_{h,j})\Delta_{l}(x)/6\right\}$
$\leq\frac{6\Delta_{l}(x)}{\sqrt{n}\mu(A_{h,j})\Delta_{l}(x)}\exp\left(-\frac{n\mu(A_{h,j})^{2}\Delta_{l}(x)^{2}/6^{2}}{4\sigma_{Z}^{2}}\right)+O\left(\frac{\sigma_{Z}}{nh_{n}^{d}}\right)$
$\leq\frac{6}{\sqrt{n}\mu(A_{h,j})}\exp\left(-\frac{n\mu(A_{h,j})^{2}\Delta(x)^{2}}{144\sigma_{Z}^{2}}\right)+O\left(\frac{\sigma_{Z}}{nh_{n}^{d}}\right).$
For $h$ small enough, this together with the arguments in (21), (5.1), (23), (5.3) and (5.3) imply
$J_{n,k,l}^{(3)}$
$\leq 2\sum_{j}{\mathbb{I}}_{\{h^{d_{a}}<\mu(A_{h,j})\}}\int_{A_{h,j}}\frac{6}{\sqrt{n}\mu(A_{h,j})}\exp\left(-\frac{n\mu(A_{h,j})^{2}\Delta(x)^{2}}{144\sigma_{Z}^{2}}\right)\mu(dx)$
$\quad+2\sum_{j}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu(A_{h,j})>0\}}\int_{A_{h,j}}\frac{6}{\sqrt{n}\mu(A_{h,j})}\exp\left(-\frac{n\mu_{a}(A_{h,j})^{2}\Delta(x)^{2}}{144\sigma_{Z}^{2}}\right)\mu_{a}(dx)$
up to an $O\left(\frac{\sigma_{Z}}{nh^{d}}\right)$ term, and henceforth
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(3)}=O\bigg{(}\bigg{(}\frac{\sigma_{Z}^{2}}{nh_{n}^{2d_{a}}}\bigg{)}^{(1+\min(\gamma,\gamma_{2}))/2}\bigg{)}+O\left(\frac{\sigma_{Z}}{nh_{n}^{d}}\right).$